# Pulse-train propagation in nonlinear Kerr media governed by higher-order dispersion

### v1

@@ Abstract
We discover three novel classes of pulse-train waveforms in an optical Kerr nonlinear medium possessing all orders of dispersion up to the fourth order. We show that both single- and double-humped pulse-trains can be formed in the nonlinear medium. A distinguishing property is that these structures have different amplitudes, widths and wavenumbers but equal velocity which depends on the three dispersion parameters. More importantly, we find that the relation between the amplitude and duration of all the newly obtained pulse-trains is determined by the sign of a joint parameter solely. The results show that those optical waves are general, in the sense that no specified conditions on the material parameters are assumed. Considering the long-wave limit, the derived pulse-trains degenerate to soliton pulses of the quartic and dipole kinds.
pacs:
05.45.Yv, 42.65.Tg
@@ I Introduction
The transmission of soliton pulses in nonlinear media has been the focus of continuing interest, stimulated by their important potential applications in long-distance optical communication systems and optical switching devises [1]. The formation of such wave packets in a monomode optical fiber results from the exact balance between the pulse-spreading effects of group-velocity dispersion and self-phase modulation effects due to the intensity-dependent refractive index of the fiber [2]. Experimental observation of solitons have been reported in various optical systems, including single-mode optical fibers [3], femtosecond lasers [4] and bulk optical materials [5].
In the picosecond regime, the propagation of light pulses inside a single-mode optical fiber was shown to be governed by the nonlinear Schrödinger equation (NLSE), which is commonly derived with use of the slowly varying envelope approximation [8]. This physically relevant model allows for either bright or dark solitons in the anomalous and normal dispersion regimes, respectively [9]. However, for much shorter light pulses, there is also appearance of additional higher-order effects in the optical system. For example, the third-order dispersion significantly influence the short pulses whose widths are nearly 50 femtosecond as they propagate through the nonlinear medium [10, 11]. The fourth-order dispersion is also found to be important when pulses are shorter than 10 femtosecond [10, 11]. In such a situation, extensions of the NLSE model that take into account the influence of diverse physical phenomena on short-pulse propagation and generation are therefore essential to describe the wave dynamics accurately. One notes that higher-order effects introduce several new important phenomena into the system dynamics that are absent in Kerr nonlinear medium.
@@
A great interest has been given recently to the generation of pulse-trains in various physical systems, including birefringent optical fibers [12, 13, 14, 15, 16], negative-index materials [17], Bose-Einstein condensates [18], etc. It can be relevant to note that such kind of pulse shape is interesting in itself as another type of physically relevant nonlinear waves in optical fibers besides the soliton-type waveform. These structures exhibit many interesting properties that make them potentially in problems of signal propagation in nonlinear fibers [19]. It is intersting to note that pulse-trains can be produced experimentally. For example, the formation of cnoidal wave has been recently demonstrated in a fiber laser either in the net normal or net anomalous cavity dispersion regime [20]. Importantly, such arrays of pulses may be created by the modulational instability of the continuous-wave signal [21], or from a dual-frequency pump [22]. In optics, such nonlinear waves play a pivotal role in the analysis of the data transmission in fiber-optic telecommunications links [23].
Identification of new kinds of pulse-trains in optical systems is an issue worth considering. It is pertinent to mention that the finding of such structures may lead to novel dynamical behaviors of waves and opens the possibility for experimental investigations of various wave phenomena. In this work, we demonstrate for the first time that an optical Kerr nonlinear medium exhibiting all orders of dispersion up to the fourth order supports the existence of three novel classes of pulse-trains which take either single- or double-hump forms.
@@
The paper is structured as follows. Section II provides the analytic framework necessary to find the traveling wave solutions of the extended NLSE model that governs the transmission of extremely short pulses throught an optical fiber medium exhibiting all orders of dispersion up to the fourth order. In Section III, we derive analytical pulse-train solutions of the model and their characteristics. We also discuss here the propagation properties of obtained pulse-trains in the fiber medium. Finally, some conclusions are given in Sec. VI.
@@ II Model and traveling waves
We consider the evolution of an optical pulse in a highly dispersive fiber medium having a Kerr nonlinearity. The governing model for the slowly varying envelope $\psi$ of the optical pulse can be described by the extended NLSE [6, 7],
@@
$i\frac{\partial\psi}{\partial z}=\alpha\frac{\partial\psi^{2}}{\partial\tau^{2}}+i\sigma\frac{\partial\psi^{3}}{\partial\tau^{3}}-\epsilon\frac{\partial\psi^{4}}{\partial\tau^{4}}-\gamma\left|\psi\right|^{2}\psi.$ (1)
@@
where $\tau=t-\beta_{1}z$ is the retarded time, $z$ is the longitudinal coordinate, $\alpha=\beta_{2}/2$, $\sigma=\beta_{3}/6$, $\epsilon=\beta_{4}/24$, and $\gamma$ is the nonlinear parameter. Here the parameters $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, and $\beta_{4}$ are the first-, second-, third-, and fourth-order dispersion coefficients, respectively.
@@
Theoretical investigations based on Eq. (1) are mainly concentrated on the existence and propagation properties of localized pulses due to their fundamental importance in the understanding of various physical phenomena in the system. For example, Kruglov and Harvey [6] demonstrated the existence of stable solitary waves with a $\mathrm{sech}^{2}$ shape for the model. Triki and Kruglov [7] discussed the dipole soliton solution of Eq. (1) in the presence the inhomogeneities of media. Cavalcanti et al. [24] analyzed the modulational instability phenomena of Eq. (1) in the region of the minimum group-velocity dispersion. Karpman et al. [25, 26] investigated the time behavior of the amplitudes, velocities, and other parameters of radiating solitons. Shagalov [27] studied the effect of the third- and fourth-order dispersions on the modulational instability. Roy et al. [29] discussed the role of third- and fourth-order dispersions in the radiation emitted by fundamental soliton pulses in the form of dispersive waves within the framework of the dimensionless form of the model (1). Kruglov [28] obtained the periodic wave solutions that take the $\mathrm{cn}^{2}$ shape for this governing equation [28]. In what follows, we present three novel exact pulse-train solutions with interesting properties for the extended NLSE (1), which are obtained without necessarily assuming a specified condition on the fiber parameters.
@@
In order to find some interesting analytical solutions of the model (1), we consider a solution of the form
$\psi(z,\tau)=\rho(\xi)\exp[i(\kappa z-\delta\tau+\theta)],$ (2)
@@
where $\rho(\xi)$ is a real function, $\xi=\tau-qz$ is a traveling coordinate with $q=v^{-1}$ being the inverse velocity. Also, $\kappa$ and $\delta$ are the respective real parameters representing the wavenumber and frequency shift, while $\theta$ represents the phase of the pulse at $z=0$.
Substitution of Eq. (1) into the extended NLSE (2) leads to a set of coupled differential equations:
$(\sigma+4\epsilon\delta)\frac{d^{3}\rho}{d\xi^{3}}+(q-2\alpha\delta-3\sigma\delta^{2}-4\epsilon\delta^{3})\frac{d\rho}{d\xi}=0,$ (3)
@@
$\epsilon\frac{d^{4}\rho}{d\xi^{4}}-(\alpha+3\sigma\delta+6\epsilon\delta^{2})\frac{d^{2}\rho}{d\xi^{2}}+\gamma\rho^{3}-(\kappa-\alpha\delta^{2}-\sigma\delta^{3}-\epsilon\delta^{4})\rho=0.$ (4)
@@
Here, we refer the solution of the extended NLSE where the function $\psi(z,\tau)$ is given by Eq. (2) with $\rho(\xi)\neq\mathrm{constant}$ as non-plain wave or traveling wave solution.
The system of Eqs. (3) and (4) with $\epsilon\neq 0$ yields the non-plain wave solutions if and only if the next relations are satisfied:
@@
Note that Eq. (3) is satisfied for an arbitrary function $\rho(\xi)$ according to the conditions in Eq. (5) with $\epsilon\neq 0$. The relations in Eq. (5) leads to the next expression for the velocity $v=1/q$ defined in the retarded frame,
@@
The latter relation shows that the dispersion parameters $\beta_{i}$ $(i=2,3,4)$ included in the parameters $\alpha$, $\sigma$ and $\epsilon$ have a major influence on the wave velocity. This implies that the velocity of the pulse can be controlled by an appropriate manipulation of the dispersion parameters.
With use of the relations in Eq. (5), the system of Eqs. (3) and (4) reduces to the ordinary nonlinear differential equation,
$\epsilon\frac{d^{4}\rho}{d\xi^{4}}+b\frac{d^{2}\rho}{d\xi^{2}}-c\rho+\gamma\rho^{3}=0,~{}~{}~{}~{}~{}~{}~{}$ (7)
where the parameters $b$ and $c$ are
@@
Equation (7) describes the evolution of the field amplitude in the highly dispersive optical fiber. Now the natural question is whether the amplitude equation (7) possesses exact pulse-train solutions without any specified condition on the coefficients or not. In following, we have identified novel families of pulse-train solutions with some interesting properties for this evolutional equation in presence of all physical parameters.
@@ III.1 Pulse-train solutions
Here we show for the first time that Eq. (7) allows for exact pulse–train solutions which are composed of the product of Jacobi elliptic functions. We find that this type of pulse shape can exist for Eq. (1) in three different classes.
Figure 1: Intensity profiles of pulse-train solutions with parameters $\gamma=2$, $k=0.9$, $\xi_{0}=0$ (a) pulse-train solution (16) when $\alpha=-1,$ $\sigma=-0.25,$ $\epsilon=-0.25$ (b) pulse-train solution (25) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5$ (c) pulse-train solution (34) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5.$
Figure 2: Intensity profiles of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 3: Evolution of pulse-train solutions with parameter $k=0.9$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 4: Evolution of pulse-train solutions with parameter $k=0.999$ (a) pulse-train solution (16) (b) pulse-train solution (25) (c) pulse-train solution (34). Other parameters are same as given in Fig. 1.
Figure 5: Dependence of the velocity of pulse-trains $v$ on the group velocity dispersion parameter $\alpha$ for $\sigma=1$ and different values of $\epsilon$, $\epsilon=-0.25$ (thick line), $\epsilon=-0.3$ (dashed line) and $\epsilon=-0.35$ (dotted line).

### v2

# Pulse-train propagation in a nonlinear Kerr medium governed by higher-order dispersion

@@ Abstract
We discover three novel classes of pulse-trains in an optical Kerr nonlinear medium possessing all orders of dispersion up to the fourth order. We show that both single- and double-humped pulse-trains can be formed in the nonlinear medium. A distinguishing property is that these structures have different amplitudes, widths and wave numbers but equal velocity which depends on the three dispersion parameters. More importantly, we find that the relation between the amplitude and duration of all the newly obtained pulse-trains is determined by the sign of a joint parameter solely. The results show that those optical waves are general, in the sense that no specified conditions on the material parameters are taken on. Considering the long-wave limit, the derived pulse-trains degenerate to soliton pulses of the quartic and dipole kinds. Results in this study may be useful for experimental realization of pulse-trains in highly dispersive optical fibers and further understanding of their optical transmission properties.
Keywords: Pulse-train, nonlinear Kerr medium, Higher-order dispersion, Quartic and dipole solitons, Nonlinear Schrödinger equation.
@@ I Introduction
The transmission of envelope solitons in waveguiding media has been the focus of continuing interest, stimulated by their important potential applications in optical communication systems and optical switching devises [1]. The formation of such wave packets in a monomode optical fiber results from an interplay between nonlinear self-phase modulation and group-velocity dispersion effects [2]. One notes that the experimental observation of solitons has been demonstrated in diverse optical systems, including nonlinear optical fibers [3], glass waveguides [4] and femtosecond ring lasers [5]. Recently, the properties of soliton pulses have been studied extensively due to their important applications for optical transmission and information processing [6].
Usually, the evolution of picosecond light pulses inside a monomode fiber has been more generally governed by the nonlinear Schrödinger equation (NLSE) that is commonly derived with use of the slowly varying envelope approximation [7]. Such a pertinent model has admitted the bright pulse soliton solution in the anomalous dispersion regime while in the normal dispersion region of the fiber it has exhibited the dark-type soliton solution [8]. However, when much shorter optical pulses have to be injected in the fiber, many higher-order dispersive and nonlinear effects have become significant and should have been taken into consideration [9]. For example, the third-order dispersion has considerably influenced the short pulses whose widths are about $50$ femtoseconds as they propagate through the nonlinear medium [10, 11]. The quartic dispersion has been also found to be significant when the light pulses are shorter than $10$ femtoseconds [10, 11]. In such a case, extensions of the cubic model taking into account the specific contributions of various higher-order effects on the light propagation have been therefore essential to describe the wave dynamics accurately [12]. It has been worth to noting that the occurrence of higher-order processes gives rise to a multitude of novel phenomena which are not found in the case of the pure Kerr media [13]. One notes that besides the NLSE and its variants, important results have been also obtained with recent studies discussing diverse nonlinear models which find applications in various sciences and engineering fields, including magnetic field [14], fluid dynamics [15, 16], magneto-optics [17], quantum mechanics [18], and plasma physics [19]. These models have been used to understand and simulate complex phenomena like solitons in magnetic materials, shallow-water waves in oceans, and the propagation of nonlinear waves in various media. We also note that significant results have been recently obtained by studying the existence and dynamics of various new soliton structures in nonlinear physical systems governed by the NLS-type equations, thus illustrating their significance in fields like optics and Bose-Einstein condensates [20, 21, 22]. Techniques like bilinear auto-Bäcklund transformations and similarity reductions have been found to be crucial for finding solutions for such models and understanding system behaviors.
@@
Nowadays, a great interest has been given to the generation of pulse-trains in various physical systems, including birefringent optical fibers [23, 24, 25, 26, 27], negative-index materials [28], Bose-Einstein condensates [29], etc. It can be relevant to note that such type of pulse shape is attracting in itself as another kind of physically relevant nonlinear waves in optical fibers besides the soliton waveforms. These structures exhibit many interesting properties that make them potentially in problems of signal propagation in nonlinear fibers [30]. It is worth mentioning that pulse-trains can be produced experimentally. For example, the formation of cnoidal waves has been recently demonstrated in a fiber laser either in the net anomalous or net normal cavity dispersion regime [31]. Importantly, such pulse-trains may be created from a dual-frequency pump [32] or by the modulational instability of the continuous-wave signal [33]. In optics, such nonlinear structures play a pivotal role in analyzing the data transmission in optical communication links [34].
Recently, a kind of extended NLSE with second-, third-, and fourth-order dispersion terms has been shown to support a stable solitary wave solution with a $\mathrm{sech}^{2}$ shape which is also called quartic soliton [35]. Such kind of stable localized pulses could find practical applications in communications, ultrafast lasers, and slow-light devices [35]. Due to its physical importance in describing femtosecond pulse propagation in highly dispersive optical fibers, such equation has been analyzed from different points of view [35, 36, 37]. With regards to the periodic waves, one of the present authors [37] has analytically solved this model and obtained results for periodic wave solutions of the $\mathrm{cn}^{2}$ type. However, analytical pulse-train solutions of this extended NLSE that are expressed in terms of products of two different Jacobi elliptic functions have not been reported yet.
Identification of new kinds of pulse-trains in optical systems is an issue worth considering. It is pertinent to mention that the finding of such structures may lead to novel dynamical behaviors of waves and opens the possibility for experimental investigations of various wave phenomena. In this study, we show for the first time that an optical Kerr nonlinear system displaying dispersion effects up to the fourth order supports the existence of three novel classes of pulse-trains which take either single- or double-hump forms. The main novelty of these solutions is that the periodic pulse-trains that are composed of products of Jacobian elliptic functions are firstly reported for the extended NLSE with third- and fourth-order dispersions, the interesting amplitude-duration relationship for these structures is found, and the propagation of pulse-train solutions is firstly represented by numerical simulations. In particular, the outcomes presented below constitute the first analytical demonstration of possible pulse-train generation in highly dispersive optical fibers, which may have potential application for the further experiments and research in nonlinear optics. These results will be helpful to nonlinear optics since this kind of nonlinear waves serves as a model of pulse train propagation in optics fibers [34] due to their structural stability with respect to the small input profile perturbations and collisions [38].
@@
The rest of paper is structured as follows. In Sec. II, we present the analytic framework necessary to find the traveling wave solutions of the extended NLSE model which governs the transmission of extremely short light pulses through a fiber system depicting higher-order dispersion effects. In Sec. III, we derive three different analytic pulse-train solutions of the governing equation and their characteristics. We also discuss here the propagation properties of obtained pulse-trains in the fiber medium. In Sec. IV, we numerically simulate the propagation of the obtained pulse-train solutions. The stability of the solutions is discussed numerically in Sec. V. The conclusions of this paper will be then summarized in Sec. VI.
@@ II Model and traveling waves
We consider the propagation of ultrashort light pulses through a nonlinear fiber medium exhibiting the effects of higher-order dispersions as well as self-phase modulation nonlinearity. The governing model for the slowly varying envelope $\psi$ of the light pulse can be described by the extended NLSE [35, 36],
@@
$\mathrm{i}\frac{\partial\psi}{\partial z}=\alpha\frac{\partial^{2}\psi}{\partial\tau^{2}}+\mathrm{i}\text{\/}\sigma\frac{\partial^{3}\psi}{\partial\tau^{3}}-\epsilon\frac{\partial^{4}\psi}{\partial\tau^{4}}-\gamma\left|\psi\right|^{2}\psi,$ (1)
@@
where $z$ is the longitudinal coordinate, $\tau=t-\beta_{1}z$ is the retarded time, $\alpha=\beta_{2}/2$, $\sigma=\beta_{3}/6$, $\epsilon=\beta_{4}/24$, while $\gamma$ is the fiber nonlinearity coefficient that results in self-phase modulation. Here $\beta_{j}$ ($j=1,2,3,4$) represent, respectively, the dispersion coefficients of the first, second, third, and fourth order.
As an important generalization, the extended NLSE (1) includes, besides the group-velocity dispersion and self-phase modulation terms that constitute the cubic NLS model, the effects of third- and fourth-order dispersions. These higher-order effects arise in highly dispersive optical systems including dispersion-shifted fibers [39], photonic crystal waveguides [40], and silicon-based waveguides [41]. It is relevant to note here that the third- and fourth-order dispersions occur in the region of minimum group-velocity dispersion [42], in which they become crucial for describing femtosecond pulse behavior.
@@
Theoretical investigations based on Eq. (1) are mainly concentrated on the existence and propagation properties of localized pulses due to their fundamental importance in the understanding of various physical phenomena in the system. As previously mentioned, Kruglov and Harvey [35] demonstrated the existence of stable solitary waves with a $\mathrm{sech}^{2}$ shape for this model. Triki and Kruglov [36] discussed the dynamics of dipole soliton waveforms of Eq. (1) in the presence the inhomogeneities of media. Kruglov [37] obtained the periodic wave solutions that take the $\mathrm{cn}^{2}$ shape for this underlying equation [37]. Cavalcanti et al. [42] analyzed the modulation instability of the model (1) in the region near the zero-dispersion wavelength. Karpman et al. [43, 44] investigated the resonant radiation and evolution of a soliton described by the model (1). Shagalov [45] studied the influence of high dispersion terms in the governing equation (1) on the modulational instability of nonlinear waves. Roy et al. [46] discussed the roles of high-order dispersions in the generation and control of dispersive waves. In what follows, we present three novel analytic pulse-train solutions with interesting properties for the extended NLSE (1), which are obtained without necessarily assuming a specified condition on the fiber parameters. Here, the ansatz method which is one of the effective and powerful techniques for finding analytic localized and periodic wave solutions of nonlinear evolution equations [14, 15, 16], is used to obtain various types of pulse-trains solutions for the studied model.
@@
In order to find some interesting analytical solutions of the model (1), we make the transformation [35, 36],
$\psi(z,\tau)=\rho(\xi)\exp[\mathrm{i}(\kappa z-\delta\tau+\theta)],$ (2)
@@
with $\rho(\xi)$ is the real amplitude and $\xi=\tau-qz$ is the traveling coordinate. Here the inverse velocity $q=v^{-1}$, the frequency shift $\delta$, the wave number $\kappa$, and the phase constant $\theta$ (at $z=0$) are all real parameters to be determined.
Inserting Eq. (2) into the extended NLSE (1), we get the differential equations:
$(\sigma+4\epsilon\delta)\frac{\mathrm{d}^{3}\rho}{\mathrm{d}\xi^{3}}+(q-2\alpha\delta-3\sigma\delta^{2}-4\epsilon\delta^{3})\frac{\mathrm{d}\rho}{\mathrm{d}\xi}=0,$ (3)
@@
$\epsilon\frac{\mathrm{d}^{4}\rho}{\mathrm{d}\xi^{4}}-(\alpha+3\sigma\delta+6\epsilon\delta^{2})\frac{\mathrm{d}^{2}\rho}{\mathrm{d}\xi^{2}}+\gamma\rho^{3}-(\kappa-\alpha\delta^{2}-\sigma\delta^{3}-\epsilon\delta^{4})\rho=0.$ (4)
@@
Here, one refers the waveform solution of the model under consideration where the envelope function $\psi(z,\tau)$ is defined by the representation (2) with $\rho(\xi)\neq\mathrm{constant}$ as traveling wave or non-plain wave solution.
Equations (3) and (4) with non-zero quartic dispersion coefficient (i.e., $\epsilon\neq 0$) give the non-plain wave solutions when the following relations are fulfilled:
@@
Notice that the differential equation (3) is fulfilled for an arbitrary amplitude function $\rho(\xi)$ in accordance with the conditions in Eq. (5) with $\epsilon\neq 0$. Additionally, Eq. (5) allows us to obtain the following expression for the wave velocity $v=1/q$,
@@
The latter relation shows that the dispersion parameters $\beta_{i}$ $(i=2,3,4)$ included in the parameters $\alpha$, $\sigma$ and $\epsilon$ have a major influence on the wave velocity. This implies that the velocity of the pulse can be efficiently controlled by varying these parameters.
Now using Eqs. (4) and (5), one can find an equation for $\rho$ as
$\epsilon\frac{\mathrm{d}^{4}\rho}{\mathrm{d}\xi^{4}}+b\frac{\mathrm{d}^{2}\rho}{\mathrm{d}\xi^{2}}-c\rho+\gamma\rho^{3}=0,$~~~~~~~ (7)
with
@@
Equation (7) describes the evolution of the wave amplitude in the highly dispersive nonlinear fiber. Now the natural question is whether the amplitude equation (7) possesses analytic pulse-train solutions without any specified condition on the coefficients or not. In following, we have identified novel families of pulse-train solutions with some interesting properties for this evolution equation taking into account the influence of all linear and nonlinear processes on pulse propagation.
@@ III.1 Pulse-train solutions
Here we show for the first time that Eq. (7) allows for analytic pulse-train solutions which are expressed in terms of the product of different elliptic functions. We find that this type of pulse shape can exist for Eq. (1) in three different classes.
Figure 1: Intensity profiles of pulse-train solutions with parameters $\gamma=2$, $k=0.9$, $\xi_{0}=0$; (a) pulse-train solution (16) when $\alpha=-1,$ $\sigma=-0.25,$ $\epsilon=-0.25$; (b) pulse-train solution (25) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5$; (c) pulse-train solution (34) when $\alpha=0.49,$ $\sigma=1,$ $\epsilon=0.5.$
Figure 2: Intensity profiles of pulse-train solutions with parameter $k=0.9$; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34). Other parameters are the same as given in Fig. 1.
Figure 3: Evolution of pulse-train solutions with parameter $k=0.999$; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34). Other parameters are the same as given in Fig. 1.
Figure 4: Evolution of pulse-train solutions with parameter $k=0.999$; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34). Other parameters are the same as given in Fig. 1.
Figure 5: Dependence of the velocity of pulse-trains $v$ on the group velocity dispersion parameter $\alpha$ for $\sigma=1$ and different values of $\epsilon$; $\epsilon=-0.25$ (thick line), $\epsilon=-0.3$ (dashed line) and $\epsilon=-0.35$ (dotted line).
Figure 6: Evolution of the intensity distribution of the extended NLSE; (a) pulse-train solution (16), (b) pulse-train solution (25), (c) pulse-train solution (34). Parameters are given in the text.